The total profit function,​ P(x), for a company producing x thousand units is given by P(x)=−2x^2+38x−120. Find the values of x for which the company makes a profit. [Hint: The company makes a profit when P(x)>0.] Explain and justify your answer.
Found 2 solutions by ewatrrr, ikleyn:
Answer by ewatrrr(24785)


Hi
  P(x)=−2x^2+38x−120 = x^2 - 19x + 60 = (x - 4)(x-15)
 P(4) = 0  and P(15) = 0     4 ≤  x  ≤ 8 will either break even or show a profit
  P(10) shows the maximum profit.



Wish You the Best in your Studies.


Answer by ikleyn(52781)

.

    P(x) = −2x^2 + 38x − 120 = -2*(x^2 - 19x + 60) = -2*(x - 4)(x-15)


    P(4) = 0  and P(15) = 0     4 ≤  x  ≤ 15   will either break even or show a profit

    
    P(9.5) shows the maximum profit.



    



Find several important differences from the post by @ewatrrr.